For each of the following equations, determine whether demand is elastic, inelastic or unitary elastic at the given price.

a) P = 50 – 0.1Q and P = Rs. 20
b) 20P =1500-Q and P=Rs.5
c) 4P =100-Q and P=Rs. 20
1
Expert's answer
2015-04-14T09:52:32-0400
Point price elasticity is calculated using the formula: Ed = P/Q*Q(P)'
a) P = 50 – 0.1Q, Q = 500 - 10P, P = Rs. 20, so Ed = 20/300*(-10) = -0.66, so demand is inelastic (|Ed| < 1).
b) 20P =1500-Q, Q = 1500 - 20P, P=Rs.5, so Ed = 5/1400*(-20) = -0.07, so demand is very inelastic (|Ed| < 1).
c) 4P =100-Q, Q = 100 - 4P, P=Rs. 20, so
Ed = 20/20*(-4) = -4, so demand is elastic (|Ed| > 1).